In 2008, Pierre’s Pizza and Burger Palace together had revenues totaling $470,000. If Burger Palace took
in $90,000 less than Pierre’s Pizza, how much did each take in as revenue?

Found 2 solutions by math_tutor2020, ikleyn:
Answer by math_tutor2020(3816)


The value 2008 is there as a distraction. It won't be used, so we can ignore it entirely.

Think of "470,000" as "470 thousand".
Same goes for 90,000
Since "thousand" shows up everywhere, I'll drop the tag and from now on all values are in thousands
Eg: something like P = 55 means "55,000" or "55 thousand".

P = Revenue for Pierre's pizza (in thousands of dollars)
B = Revenue for Burger palace (in thousands of dollars)

Since their revenues total 470 thousand, we can say
P+B = 470

Furthermore, the Burger Palace took in 90 thousand less than Pierre's Pizza, so B = P - 90

Let's apply substitution to solve
P+B = 470
P+P-90 = 470 ... replace B with P-90
2P-90 = 470
2P = 470+90
2P = 560
P = 560/2
P = 280
Recall that all values are in thousands, so P = 280 represents 280,000

Use this to find B
B = P - 90
B = 280 - 90
B = 190
This represents 190,000

----------------------------------------------

Answers:

Pierre's Pizza had revenue of $280,000
Burger Palace had revenue of $190,000

Answer by ikleyn(52756)

.

Write an equation as you read the problem

    P + B = 470000,


and then

    x + (x-90000) = 470000


(x stands for P = Pierre).


Continue

    2x = 470,000 + 90,000 = 560,000

     x                    = 560000/2 = 280000.


ANSWER.  $280,000 for P  and  $280,000 - $90,000 = $190,000 for B.

Solved.

Is everything clear to you in my post ?